# LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces

### v2

@@ 4.1 Experimental Setup
4.2 Results
We benchmark our model against various domain adaptation methods that have reported results on the considered datasets [77, 68, 25, 75, 7, 73]. Besides classic UDA methods, we also compare with recent approaches using language guidance [27]. Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, while for Ego2Exo, we use per-class mean accuracy as reported in [27].
DomainNet. Table 1 presents the results on the DomainNet dataset, encompassing $12$ domain adaptation scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P). LAGUNA is compared against previous state-of-the-art methods utilizing the Swin-B backbone to ensure a fair comparison. Notably, LAGUNA surpasses all prior work in all $12$ scenarios, achieving an average accuracy improvement of $+3.32\%$.
[…]
These consistent improvements across all tested datasets, with a total of 18 scenarios, underscore the effectiveness of our approach both when considering classic UDA algorithms and recent methods exploring the use of language guidance to assist domain adaptation.
@@ 4.3 Ablation Study
In this section, we analyze the effect of the individual components of LAGUNA to provide an in-depth assessment of model performance and justify our design choices. Specifically, we investigate (a) the impact of our methodological elements, (b) the effects of different reference structures, (c) the behavior of LAGUNA with different ratios of pseudo-labels, and (d) the qualitative illustration of the alignment of relative and absolute spaces. More ablations are reported in the Supplementary Material.
Methodological Elements. In Table 4, we ablate the main methodological elements of LAGUNA on GeoImnet. We set up this ablation by removing the core elements of LAGUNA, namely the structure loss ($\mathcal{L}_{S}$ Eq. (6)), the use of domain-agnostic reference anchors $\mathcal{A}$, the learnable anchors $\mathcal{A}_{t/s}$, the cross-domain attention (CD Attn. - Eq. (9), (10)), and the regularization loss ($\mathcal{L}_{Reg}$ Eq. (12)). We then progressively add the considered elements until we reach LAGUNA (settings (1)-(5)). For all settings, we report the relative improvement (w.r.t. previous row) and the absolute one (w.r.t. first row). In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran [27]. In setting (2), we add the reference anchors $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity. This yields an improvement of $+0.87$ with respect to (1), suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance but that absolute alignment is too restrictive. In setting (3), we extend (2) with a set of domain-independent learnable anchors ($\checkmark$*✓ ∗ in Table 4, indicating that the weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the structure-preserving loss $\mathcal{L}_{S}$ (Eq. 6). The introduction of learnable anchors enables the relative alignment of the visual domain with the reference language domain, leading to an improvement of $+1.32$ with respect to (2) and $+2.10$ over (1), which highlights the benefits of the proposed relative alignment over an absolute one. Considering domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of $+1.77$ with respect to (3) and a robust $+3.87$ with respect to the baseline (1). Adding cross-domain attention in setting (5) allows to bridge the gap between target and source representations, leading to an additional improvement of $+0.44$ with respect to (4) and an overall improvement of $+4.31$ with respect to (1). We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of $+1.16$ with respect to (5) and an overall gain of $+5.47$ with respect to the baseline (1). Together, these improvements show the advantages of the careful design of LAGUNA, highlighting the benefits of learning domain-specific visual representations still aligned to a reference language structure.
@@
Ablation on reference structures. Fig. 3 delves into the importance of defining a semantic structure for our method (Stage 1). We ablate three models: SentenceTransformer[55], CLIP [54], and BERT [59], each with varying capabilities of extracting semantic relationships. The figure shows semantic similarity maps for $100$ randomly selected classes from GeoImnet and the average accuracy obtained by LAGUNA on the entire dataset when using the specified model. As can be seen, SentenceTransformer keeps low similarities for distinct classes (low values off-diagonal), BERT maintains high similarities for most classes but also presents some spurious correlations between distinct classes, while the behavior of CLIP is in-between. SentenceTransformer, which can better differentiate between semantic relations, leads to the best accuracy.
Figure 3: Ablation on three language models, SentenceTransformer, CLIP, and BERT, used to define the reference structure. We visualize the similarity maps of $100$ randomly selected classes from GeoImnet (green to yellow areas indicate higher similarity) and the average accuracies on all classes achieved by using each model to define the reference space. The best result is in bold, and the second best is underlined.
@@
Ablation on the ratio of target samples. LAGUNA demonstrates strong generalization capabilities even with limited target domain data, as shown in Table 5. We progressively increase the amount of pseudo-labeled target data used for training (10%, 20%, 30%, 50%, 75%) and observe the impact on performance. Notably, LAGUNA achieves high accuracy even with only $10\%$ of the target samples and attains state-of-the-art results with just $20\%$. This efficiency underscores the benefits of our structure-driven approach, which organizes sample projections within domain-specific representation spaces while leveraging a reference geometrical structure to guide their placement.
Dataset 10% 20% 30% 50% 75% 100%
GeoImnet 61.40 63.43 64.21 66.42 66.71 68.68
GeoPlaces 55.35 57.85 58.95 60.02 60.75 62.33
Avg. 58.38 60.64 61.58 63.22 63.72 65.40
Table 5: Ablation using different target data ratios. We report average accuracy on GeoImnet and GeoPlaces. The result in Green indicates the ratio needed for LAGUNA to obtain SOTA accuracy.
Figure 4: Visualization of Source and Target (Cross-Domain) visual features of 1000 randomly selected samples from GeoImnet with t-SNE. We illustrate their position in latent space from two perspectives: left) absolute coordinates and right) relative coordinates. Notably, in absolute coordinates, the two representation spaces are quite distinct, adhering to our training strategy without direct alignment. Relative encodings w.r.t $\mathcal{A}_{s}$ and $\mathcal{A}_{t}$ are aligned.
@@
Qualitative illustration of embedding spaces. LAGUNA uniquely aligns representation structures without enforcing alignment in absolute coordinates but rather leveraging a reference structure defined by a language model and the class labels to guide the alignment of inter-class affinities. Figure 4 provides a qualitative validation of this approach. Using t-SNE, we visualize the projection of $1000$ randomly selected samples from the GeoImnet validation set, where the source and the target domains belong to the USA and Asia, respectively. Notably, the features are well separated in absolute coordinates, reflecting their unique domain-related visual properties. However, when using relative encodings with domain-specific learnable anchors, the features blend together due to the structure-preserving loss $\mathcal{L}_{S}$. This demonstrates that semantic alignment can be achieved without forcing representations into a shared space, allowing each domain to preserve its unique characteristics while adhering to a common structure.
@@ 5 Conclusions
This work introduces LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation. LAGUNA defines a reference representation space structure based on domain-agnostic class semantic similarities encoded by a language model, ensuring the organization of sample projections reflects this structure while preserving domain-specific characteristics. By conditioning the classifier to adhere to this structure, LAGUNA encourages structural similarity across domain-specific latent spaces, retaining unique features for improved classification. This is achieved through pseudo-labeling and learnable domain-specific anchors, guided by a loss function that prioritizes mimicking geometrical associations over direct representation alignment. Extensive experiments demonstrate LAGUNA’s superior performance compared to existing methods. This research highlights the importance of structural alignment and language-guided learning in domain adaptation. Future work could explore extending LAGUNA to more complex scenarios and multi-modal data.

### v3

@@ 4.3 Ablation Study
We analyze the effect of the individual components of LAGUNA to provide an in-depth assessment of model performance and justify design choices. We investigate (a) the impact of our methodological elements, (b) the effects of different reference structures, (c) the behavior of LAGUNA with different ratios of pseudo-labels, and (d) the qualitative illustration of the alignment of relative and absolute spaces. More ablations are reported in the Supplementary Material.
@@
Methodological Elements. Table 4, ablates the main methodological elements of LAGUNA on GeoImnet. We set up this ablation by removing LAGUNA’s core elements, namely the structure loss ($\mathcal{L}_{S}$ Eq. (6)), domain-agnostic reference anchors $\mathcal{A}$, learnable anchors $\mathcal{A}_{t/s}$, cross-domain attention (CD Attn. - Eq. (9), (10)), and the regularization loss ($\mathcal{L}_{Reg}$ Eq. (12)). We then progressively add these elements (settings (1)-(5)) until we reach LAGUNA. For all settings, we report the relative improvement (w.r.t. previous row) and the absolute one (w.r.t. first row).
Absolute alignment In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran [27]. In setting (2), we add the reference anchors $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity. This yields an improvement of $+0.87$ with respect to (1), suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance and that absolute alignment is too restrictive.
Relative alignment In setting (3), we extend (2) with a set of domain-independent learnable anchors ($\checkmark$*✓ ∗ in Table 4, indicating that the weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the structure-preserving loss $\mathcal{L}_{S}$ (Eq. 6). The introduction of learnable anchors enables the relative alignment of the visual domain with the reference language domain, leading to an improvement of $+1.32$ with respect to (2) and $+2.10$ over (1), which highlights the benefits of the proposed relative alignment over an absolute one.
Domain-Specific Anchors Adding domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of $+1.77$ with respect to (3) and a robust $+3.87$ with respect to the baseline (1).
Cross-Domain Attention Adding cross-domain attention in setting (5) bridges the gap between target and source representations, leading to a of $+0.44$ with respect to (4) and an overall improvement of $+4.31$ with respect to (1).
Regularization Loss We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of $+1.16$ with respect to (5) and an overall gain of $+5.47$ with respect to (1). Together, these improvements show the effectiveness of LAGUNA, highlighting the benefits of learning domain-specific visual representations aligned to a reference language structure. Language models for reference structures. Fig. 3 reports semantic similarity maps for $100$ randomly selected classes form GeoImnet when three different language models are used in stage 1 (SentenceTransformer [57], CLIP [56], and BERT [61]), together with the average accuracy obtained keeping stages 2-3 fixed. Notably, accuracy follows the ability of models to distinguish distinct classes, with low off-diagonal values indicating more discriminative models. While we achieve best results with SentenceTransformer, LAGUNA’s performance is robust to different reference structures, maintaining SOTA results also when the same language model is used across all stages (i.e., BERT).
Figure 3: Similarity maps of $100$ randomly selected classes from GeoImnet (yellow for high similarity) and average accuracies.
@@
Ratio of target samples. LAGUNA demonstrates strong generalization capabilities even with limited target domain data (Table 5). We progressively increase the number of randomly selected pseudo-labeled target data used for training (10%-75%) and observe the impact on performance. Notably, LAGUNA achieves high accuracy even with only $10\%$ of the samples and attains SOTA results with just $20\%$ (compared with  [27]).
Dataset 10% 20% 30% 50% 75% 100% [27]
GeoImnet 61.40 63.43 64.21 66.42 66.71 68.68 63.92
GeoPlaces 55.35 57.85 58.95 60.02 60.75 62.33 56.58
Avg. 58.38 60.64 61.58 63.22 63.72 65.40 60.24
Table 5: Average accuracy with different percentages of target data compared with previous SOTA LaGTran [27]. LAGUNA to obtains SOTA accuracy with $20\%$ (in green).
@@
Alignment in Absolute and Realtive Coordinates. Figure 4 shows t-SNE plots and related Maximum Mean Discrepancy (MMD) scores for $1000$ randomly selected samples from the GeoImnet validation set, where source is USA and target is Asia. Notably, features are separated in absolute coordinates, while they blend together (low MMD value) in relative coordinates thanks to the structure-preserving loss $\mathcal{L}_{S}$ and domain-specific learnable anchors. This highlights that semantic alignment can be achieved without forcing representations into a shared space, allowing each domain to preserve its unique characteristics while adhering to a common structure.
Figure 4: t-SNE plots for 1000 randomly selected Source and Target samples from GeoImnet, with respective MMD [28] scores. Relative coordinates (right) are better aligned than absolute (left).
Figure 5: Ablation with BLIP-1 and BLIP-2 generated captions.
[…]
Computational Overhead. While LAGUNA involves three stages and different components, its training is quite efficient: Stage 1 needs no training, and embedding 600 class labels takes 1-2 minutes ($\sim$4GB GPU memory). On GeoImNet, with a V100, Stage 2 requires only 9 minutes per epoch, while Stage 3 adds $<$2 minutes per epoch, which is minor considering the baseline’s 65 minutes per epoch. At test time, LAGUNA only requires the visual encoder, the learned anchors, the light cross-attention layer and the final MLP classifier, bringing no noticeable overhead.
@@ 5 Conclusions
We introduced LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation. LAGUNA defines a reference representation space structure based on domain-agnostic class semantic similarities encoded by a language model, ensuring the organization of sample projections reflects this structure while preserving domain-specific characteristics. By conditioning the classifier to adhere to this structure, LAGUNA encourages structural similarity across domain-specific latent spaces, retaining unique features for improved classification. This is achieved through pseudo-labeling and learnable domain-specific anchors, guided by a loss function that prioritizes mimicking geometrical associations over direct representation alignment. Extensive experiments demonstrate LAGUNA’s superior performance compared to existing methods.